OpenStudy (anonymous):

Corey needs 24 boards that are 47 1/2 inches long. How many feet of boards should he but ? Explain.

OpenStudy (anonymous):

If each board is 47.5 in. long, then the total length of 24 boards is 24*47.5 inches. Next, you convert the inches to feet by dividing by 12 (as there are 12 inches in a foot). There you have your answer.

OpenStudy (31356):

(24 * 47.5) / 12 = _______ feet 24/2 = _______ number of boards

OpenStudy (31356):

1. (24 * 47.5) / 12 = ? 2. 8 * 12 = 96 inches 96 / 47.5 = 2.02 He can get two 47.50-inch lengths out of each 8-foot board. So he needs 12-foot boards.

OpenStudy (31356):

Number two is just extra